OpenStudy (anonymous):

Show all work to simplify 2/x - 2/(x-1) + 2/(x-2). Use complete sentences to explain how to simplify this expression. Remember to list all restrictions.

OpenStudy (anonymous):

\[\frac{ 2 }{ x }-\frac{ 2 }{ x-1 }+\frac{ 2 }{ x-1 }\]